you want to install a 1yd wide walk around a circular swimming pool. the diameter of the pool is 20yd. what is the area of the walk?
Answer by richwmiller(17219)

a=pi*r^2=pi*10^2
a=pi*r^2=pi*11^2
121*pi-100*pi=21*pi
The walkway is 21*pi yd^2